# citations.bib

### v1

@@ A.4 Fine-grained bounds for specific probability distributions
Now, we map $i$ to $(h,l)$ for appropriate integers $h\in[d-1]$ and $l\in[\binom{d}{h}]$, so that the subset $S_{i}\subset[d]$ is of size $h$ (according to the chosen ordering of subsets). Then, writing $k(h)=k(S_{i})$, we have $\bm{Z}^{\mathsf{T}}\sqrt{\bm{W}}\bm{e}_{h,l}=\sqrt{k(h)}\bm{Z}^{\mathsf{T}}\bm{e}_{h,l}=\sqrt{k(h)}\bm{z}_{h,l}$, where $\bm{z}_{h,l}$ is a $d$-dimensional vector with $1$ at entry $j$ if $j\in S_{h,l}$ and 00 otherwise. Substituting this in Eq. A.49, using $\bm{P}=\bm{I}-(1/d)\bm{1}\bm{1}^{\mathsf{T}}$ and $|S_{h,l}|=\left\lVert\bm{z}_{h,l}\right\rVert_{1}=\left\lVert\bm{z}_{h,l}\right\rVert^{2}=h$, we obtain
@@
1. It can be verified that $\sum_{i=1}^{2^{d}-1}\left\lVert\bm{u}_{i}\right\rVert^{2}=\left\lVert\bm{U}_{\mathsf{F}}\right\rVert^{2}=d-1$. The result follow from the definition of $p_{h,l}$, and $\gamma$, $\eta$ in Eq. 2.3.
2. Noting that $k(S)$ depends only on the size of the subset $S\subseteq[d]$, $p_{h,l}$ in Item 2 is obtained by direct calculation. Observe that
@@
for $h\in[d-1]$ and $l\in[\binom{d}{h}]$. Since $h(d-h)\leq d^{2}/4$ for $h\in[d-1]$, Item 2 follows from Eq. 2.3 by direct substitution.
3. We obtain $p_{h,l}$ in Item 3 by direct substitution. Since
@@
for $h\in[d-1]$ and $l\in[\binom{d}{h}]$, we obtain Item 3. ∎
@@
The sum over $1/h$ in Item 2 and over $1/\sqrt{h(d-h)}$ in Item 3 only mildly depend on $d$. Indeed,
@@
[…]
@@
Theorem A.6 allows us to directly compare the values of $\gamma$ for the different sampling strategies we consider.
@@
Denote $\gamma_{\ell_{2}^{2}}$, $\gamma_{\textnormal{ker}}$, $\gamma_{\textnormal{m-}\ell_{2}}$ to be the expressions for $\gamma$ for $\ell_{2}$-squared Item 1, kernel Item 2, and modified $\ell_{2}$ Item 3 sampling schemes respectively. Then, for all $\bm{z}\in\mathbb{R}^{2^{d}-2}$, we have
@@
Then, Eq. A.58, Eq. A.59, and Eq. A.60 follow from Theorem A.6 and Remark A.7. ∎
These bounds suggest that kernel weights perform at most a log factor worse than leverage scores, while it can perform nearly $d$ better than leverage scores. On the other hand, the performance of modified $\ell_{2}$ weights is never worse than leverage scores (up to constant factors), but can nearly do $\sqrt{d}$ better than leverage scores. In Appendix E, we explicitly construct a toy model that demonstrates such an advantage. While these results are only upper bounds on the sample complexities, we also observe similar results in experiments. Using Theorem A.6, we can derive the values of $\gamma$ and $\eta$ listed in Table 1 for the different sampling strategies as follows.
@@
This follows from Theorem A.6, Remark A.7 and the definition of $\bm{H}$. ∎
@@
The distributions considered in Theorem A.6 are actually a special case of a family of distributions, obtained by interpolating between kernel weights and leverage scores. Specifically, given $\tau\in[0,1]$, we can consider the weighted geometric mean $(k(h))^{\tau}(\lVert\bm{u}_{h,l}\rVert^{2})^{(1-\tau)}$ of $k(h)$ and $\lVert\bm{u}_{h,l}\rVert^{2}$ for $h\in[d-1]$ and $l\in[\binom{d}{h}]$. This gives rise to the distribution
@@
For $\tau=0$, we get the leverage scores (or $\ell_{2}$ squared distribution), $\tau=1$ gives the kernel weight distribution, and $\tau=1/2$ gives the modified $\ell_{2}$ distribution.
@@
$\eta_{\tau}=\begin{cases}\left(\frac{d^{2}}{4}\right)^{\tau}\mathcal{N}_{\tau}&\textnormal{ if $d$ is even}\\ \left(\frac{d^{2}-1}{4}\right)^{\tau}\mathcal{N}_{\tau}&\textnormal{ if $d$ isodd},\end{cases}$ (A.67)
@@
Using similar arguments as in Remark A.7, we can show that
@@
for $10\leq\tau<$ 1, where $\operatorname{B}_{z}(a,b)=\int_{0}^{z}t^{a-1}(1-t)^{b-1}\mathrm{d}t$ is the incomplete beta function.
@@
for $\tau=1$. For $10\leq\tau<$ 1, we do no worse than leverage score sampling. We remark that because the $\Theta$ notation hides constants, for a given dimension, one can choose an appropriate $\tau$ that minimizes these constants. It remains to see how such a strategy performs in practice.
@@ A.5 Theoretical guarantees for sampling without replacement
In this section, we prove guarantees for matrix vector multiplication estimator and the regression estimator when the rows/columns are sampled without replacement. [^5] We note that the term “sampling without replacement” is perhaps a bit of a misnomer for this type of sampling scheme. Nevertheless, we use it in order to maintain consistency with [musco2025provably]. We follow the strategy of [musco2025provably] for sampling indices without replacement.
@@
A.5.1 Approximate Matrix Multiplication
We derive the following guarantee for the matrix multiplication estimator for sampling without replacement. Since the number of samples are not fixed, we instead calculate the estimation error for a fixed expected number of samples (which is determined by the probabilities $q_{1},\dotsc,q_{r}$).
Theorem A.11 (Matrix multiplication, sampling without replacement).
@@ A.5.1 Approximate Matrix Multiplication
$\textnormal{var}(\widehat{\bm{X}})$ $=\sum_{i=1}^{r}\textnormal{var}\left(Y_{i}\frac{\bm{u}_{i}z_{i}}{q_{i}}\right)$ (A.75)
@@
Setting the right-hand-side of the above inequality equal to $\delta$ and solving for $\varepsilon$ gives us Eq. A.73. ∎
We can use the above result to derive the error bounds in terms of the function $\gamma(\bm{z})$ defined in Eq. 2.3.
@@
where we use the fact that $q_{i}=cp_{i}$ when $q_{i}<1$ in the second line, the fact that the terms are non-negative in the third line, and the definition of $\gamma$ (see Eq. 2.3) and $c\geq m_{0}$ in the third line. The result then follows from Theorem A.11. ∎
@@ A.5.2 Subspace Embedding
The result then follows from A.2. ∎
@@ A.5.3 Sketched Regression
We now combine approximate matrix multiplication guarantee (Corollary A.12) and subspace embedding guarantee (Theorem A.14) to obtain guarantee for the sketched regression estimator constructed by sampling without replacement.
@@
We closely follow the proof of Theorem A.4. Since $\bm{y}^{*}$ is the solution of $\min_{\bm{y}}\left\lVert\bm{U}\bm{y}-\bm{b}\right\rVert$, $\bm{U}\bm{y}^{*}-\bm{b}$ lies in the orthogonal complement of the range of $\bm{U}$, and therefore, $\bm{U}^{\mathsf{T}}(\bm{U}\bm{y}^{*}-\bm{b})=0$. Then, taking $\bm{z}=\bm{U}\bm{y}^{*}-\bm{b}=(\bm{I}-\bm{U}\bm{U}^{\mathsf{T}})\bm{b}$ in Corollary A.12, we can infer that using
@@
we have with probability exceeding $21-\delta/$ 2,
@@
then, with probability exceeding $21-\delta/$ 2,
@@
The remainder of the proof is the same as that of Theorem A.4. ∎
@@ Appendix B Description of past estimators
In this section we provide more details on how several existing estimators fit into the unified framework described in Section 2.
@@ B.1 KernelSHAP
KernelSHAP makes use of a subsampled and reweighted version of the constrained regression formulation Eq. 1.2 of the Shapley values. Specifically, denoting $\bm{Z}_{S}$ to be the $S$-th row of $\bm{Z}$, observe that
@@
The KernelSHAP estimator [lundberg2017unified] is then defined as
@@
As noted by [musco2025provably], this can be viewed as a constrained sketched regression problem
$\bm{\phi}^{\textup{KS}}=\operatornamewithlimits{argmin}_{\begin{subarray}{c}\bm{\phi}\in\mathbb{R}^{d}\\ \bm{1}^{\mathsf{T}}\bm{\phi}=v([d])-v(\emptyset)\end{subarray}}\left\lVert\bm{S}(\bm{Z}^{\prime}\bm{\phi}-\bm{b})\right\rVert$ (B.3)
Performing the same change of variables as in the proof of Theorem 2.1 we find that
$\bm{\phi}^{\textup{KS}}$ $=\bm{Q}\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\left\lVert\bm{S}(\bm{Z}^{\prime}(\bm{Q}\bm{x}+\alpha\bm{1})-\bm{b})\right\rVert+\alpha\bm{1}$ (B.4)
$=\bm{Q}\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\left\lVert\bm{S}(\bm{Z}^{\prime}\bm{Q}\bm{x}-(\bm{b}-\alpha\bm{Z}^{\prime}\bm{1}))\right\rVert+\alpha\bm{1}$ (B.5)
$=\bm{Q}\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\left\lVert\bm{S}(\bm{U}\bm{x}-\bm{b}_{\alpha})\right\rVert+\alpha\bm{1}.$ (B.6)
@@ B.2 Unbiased KernelSHAP
In [covert2020improving], the authors observe that the Shapley values can be expressed as
@@
Since $[\bm{Q},d^{-1/2}\bm{1}]$ form an orthonormal basis for $\mathbb{R}^{d}$,
@@ B.3 LeverageSHAP
In [musco2025provably], the authors show the typical formulation of the Shapley values Eq. 1.2 can be rewritten as an unconstrained problem
@@ Appendix C Equivalence between Lagrangian and Change of Variable Framework
$\bm{\phi}^{\textup{R}}=\operatornamewithlimits{argmin}_{\begin{subarray}{c}\bm{\phi}\in\mathbb{R}^{d}\\ \bm{1}^{\mathsf{T}}\bm{\phi}=\alpha\end{subarray}}\left\lVert\bm{C}\bm{\phi}-\bm{y}\right\rVert.$ (C.1)
@@
$\bm{\phi}^{u}=\operatornamewithlimits{argmin}_{\begin{subarray}{c}\bm{\phi}\in\mathbb{R}^{d}\end{subarray}}\left\lVert\bm{C}\bm{\phi}-\bm{y}\right\rVert=\bm{M}^{+}\bm{g}$ (C.2)
@@
In order to solve Eq. C.1, the Lagrangian method writes,
@@
As discussed in Section A.3, an alternative method to solve the constrained least squares is using the change of variable to explicitly enforce the constraint. Specifically, we re-parameterize $\bm{\phi}$ as,
@@
$\operatornamewithlimits{argmin}_{\begin{subarray}{c}\bm{x}\end{subarray}}\left\lVert\bm{C}(\alpha\bm{1}+\bm{Q}\bm{x})-\bm{y}\right\rVert=\operatornamewithlimits{argmin}_{\begin{subarray}{c}\bm{x}\end{subarray}}\left\lVert\bm{C}\bm{Q}\bm{x}-(\bm{y}-\alpha\bm{C1})\right\rVert$ (C.6)
@@
The second term in Eq. C.8 can be seen as a projection of the vector $\bm{\phi}^{u}-\alpha\bm{1}$ into the span of $\bm{Q}$ (or alternatively on the null space of $\bm{1}^{\mathsf{T}}$) with the projection matrix,
@@
Thus, plugging in this Eq. C.11 results in
@@
thus recovering Eq. C.4 by noting that $\bm{1}^{\mathsf{T}}\bm{1}=d$.
Appendix D Variance Analysis
D.1 Variance reduction
If we have a good guess what the value function $v(s)$ looks like ahead of time, then we might hope to be able to incorporate this information into our estimators.
Observe that, for any fixed vector $\bm{f}$,
$\bm{\phi}^{*}=\bm{Q}\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\left\lVert\bm{U}\bm{x}-\bm{b}_{0}\right\rVert+\alpha\bm{1}=\bm{Q}\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\left\lVert\bm{U}\bm{x}-(\bm{b}_{0}-\bm{f})\right\rVert+\alpha\bm{1}+\bm{Q}\bm{U}^{\mathsf{T}}\bm{f}$ (D.1)
and
$\bm{\phi}^{*}=\bm{Q}\bm{U}^{\mathsf{T}}\bm{b}+\alpha\bm{1}=\bm{Q}\bm{U}^{\mathsf{T}}(\bm{b}_{0}-\bm{f})+\alpha\bm{1}+\bm{Q}\bm{U}^{\mathsf{T}}\bm{f}.$ (D.2)
We can then define estimators [^8] For the special choice $\bm{f}=\sqrt{d/(d-1)}\lambda\bm{Z}^{\prime}\bm{1}$ we have that $\bm{\phi}_{\bm{f}}^{\textup{R-vr}}=\bm{\phi}_{\lambda}^{\textup{R}}$ and $\bm{\phi}_{\bm{f}}^{\textup{M-vr}}=\bm{\phi}_{\lambda}^{\textup{M}}$, and hence these estimators generalize those described in Section 2.1.
$\bm{\phi}_{\bm{f}}^{\textup{R-vr}}=\bm{Q}\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\left\lVert\bm{S}(\bm{U}\bm{x}-(\bm{b}_{0}-\bm{f}))\right\rVert+\alpha\bm{1}+\bm{Q}\bm{U}^{\mathsf{T}}\bm{f}$ (D.3)
and
$\bm{\phi}_{\bm{f}}^{\textup{M-vr}}=\bm{Q}\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}(\bm{b}_{0}-\bm{f})+\alpha\bm{1}+\bm{Q}\bm{U}^{\mathsf{T}}\bm{f}.$ (D.4)
They key observation is that, if $\bm{f}\approx\bm{b}_{0}$, then we expect the variance of $\bm{\phi}_{\bm{f}}^{\textup{R-vr}}$ and $\bm{\phi}_{\bm{f}}^{\textup{M-vr}}$ to be reduced.
We can derive theoretical guarantees. Following the same approach as Theorem 2.2, we see that
$m=O\left(\frac{\gamma(\bm{P}_{\bm{U}}(\bm{b}_{0}-\bm{f}))}{\delta\varepsilon^{2}}+\eta\log\left(\frac{d}{\delta}\right)\right)\quad\Longrightarrow\quad\mathbb{P}\left[\left\lVert\bm{\phi}^{*}-\bm{\phi}_{\bm{f}}^{\textup{R-vr}}\right\rVert<\varepsilon\right]>1-\delta,$ (D.5)
and
$m=O\left(\frac{\gamma(\bm{b}_{0}-\bm{f})}{\delta\varepsilon^{2}}\right)\quad\Longrightarrow\quad\mathbb{P}\left[\left\lVert\bm{\phi}^{*}-\bm{\phi}_{\bm{f}}^{\textup{M-vr}}\right\rVert<\varepsilon\right]>1-\delta.$ (D.6)
D.2 Ratio of mean squared errors
In Section A.2, we saw that $\gamma(\bm{z})$ and $\eta$ (see Eq. 2.3) give upper bounds on the sample complexity of matrix multiplication and regression estimators for sampling with replacement. In this section, we study the ratio of mean squared errors for different sampling strategies these estimators in the finite-sample/asymptotic regime. We find that this ratio is determined by $\gamma$ for both these estimators, as summarized below.

### v2

@@ A.4 Fine-grained bounds for specific probability distributions
Now, we map $i$ to $(h,l)$ for appropriate integers $h\in[d-1]$ and $l\in[\binom{d}{h}$], so that the subset $S_{i}\subset[d]$ is of size $h$ (according to the chosen ordering of subsets). Then, writing $k(h)=k(S_{i})$, we have $\bm{Z}^{\mathsf{T}}\sqrt{\bm{W}}\bm{e}_{h,l}=\sqrt{k(h)}\bm{Z}^{\mathsf{T}}\bm{e}_{h,l}=\sqrt{k(h)}\bm{z}_{h,l}$, where $\bm{z}_{h,l}$ is a $d$-dimensional vector with $1$ at entry $j$ if $j\in S_{h,l}$ and 0 otherwise. Substituting this in Eq.˜A.51, using $\bm{P}=\bm{I}-(1/d)\bm{1}\bm{1}^{\mathsf{T}}$ and $|S_{h,l}|=\left\lVert\bm{z}_{h,l}\right\rVert_{1}=\left\lVert\bm{z}_{h,l}\right\rVert^{2}=h$, we obtain
@@
1. It can be verified that $\sum_{i=1}^{2^{d}-1}\left\lVert\bm{u}_{i}\right\rVert^{2}=\left\lVert\bm{U}_{\mathsf{F}}\right\rVert^{2}=d-1$. The result follow from the definition of $p_{h,l}$, and $\gamma$, $\eta$ in Eq.˜2.3.
2. Noting that $k(S)$ depends only on the size of the subset $S\subseteq[d]$, $p_{h,l}$ in Item˜2 is obtained by direct calculation. Observe that
@@
for $h\in[d-1]$ and $l\in[\binom{d}{h}$]. Since $h(d-h)\leq d^{2}/4$ for $h\in[d-1]$, Item˜2 follows from Eq.˜2.3 by direct substitution.
3. We obtain $p_{h,l}$ in Item˜3 by direct substitution. Since
@@
for $h\in[d-1]$ and $l\in[\binom{d}{h}$], we obtain Item˜3. ∎
@@
The sum over $1/h$ in Item˜2 and over $1/\sqrt{h(d-h)}$ in Item˜3 only mildly depend on $d$. Indeed,
@@
$\leq 2\left(\frac{1}{\sqrt{d-1}}+\int_{1}^{d/2}\frac{1}{\sqrt{x(d-x)}}dx\right)$
@@
Theorem˜A.7 allows us to directly compare the values of $\gamma$ for the different sampling strategies we consider.
@@
Denote $\gamma_{\ell_{2}^{2}}$, $\gamma_{\textnormal{ker}}$, $\gamma_{\textnormal{m-}\ell_{2}}$ to be the expressions for $\gamma$ for $\ell_{2}$-squared Item˜1, kernel Item˜2, and modified $\ell_{2}$ Item˜3 sampling schemes respectively. Then, for all $\bm{z}\in\mathbb{R}^{2^{d}-2}$, we have
@@
Then, Eq.˜A.60, Eq.˜A.61, and Eq.˜A.62 follow from Theorem˜A.7 and Remark˜A.8. ∎
These bounds suggest that kernel weights perform at most a log factor worse than leverage scores, while it can perform nearly $d$ better than leverage scores. On the other hand, the performance of modified $\ell_{2}$ weights is never worse than leverage scores (up to constant factors), but can nearly do $\sqrt{d}$ better than leverage scores. In Appendix˜E, we explicitly construct a toy model that demonstrates such an advantage. While these results are only upper bounds on the sample complexities, we also observe similar results in experiments. Using Theorem˜A.7, we can derive the values of $\gamma$ and $\eta$ listed in Table˜1 for the different sampling strategies as follows.
@@
This follows from Theorem˜A.7, Remark˜A.8 and the definition of $\bm{H}$. ∎
@@
The distributions considered in Theorem˜A.7 are actually a special case of a family of distributions, obtained by interpolating between kernel weights and leverage scores. Specifically, given $\tau\in[0,1]$, we can consider the weighted geometric mean $(k(h))^{\tau}(\lVert\bm{u}_{h,l}\rVert^{2})^{(1-\tau)}$ of $k(h)$ and $\lVert\bm{u}_{h,l}\rVert^{2}$ for $h\in[d-1]$ and $l\in[\binom{d}{h}$]. This gives rise to the distribution
@@
For $\tau=0$, we get the leverage scores (or $\ell_{2}$-squared distribution), $\tau=1$ gives the kernel weight distribution, and $\tau=1/2$ gives the modified $\ell_{2}$ distribution.
@@
[…]
@@
Using similar arguments as in Remark˜A.8, we can show that
@@
for $0\leq\tau<1$, where $\operatorname{B}_{z}(a,b)=\int_{0}^{z}t^{a-1}(1-t)^{b-1}\mathrm{d}t$ is the incomplete beta function.
@@
for $\tau=1$. For $0\leq\tau<1$, we do no worse than leverage score sampling. We remark that because the $\Theta$ notation hides constants, for a given dimension, one can choose an appropriate $\tau$ that minimizes these constants. It remains to see how such a strategy performs in practice.
@@ A.5 Theoretical guarantees for sampling without replacement
In this section, we prove guarantees for matrix vector multiplication estimator and the regression estimator when the rows/columns are sampled without replacement. [^5] We note that the term “sampling without replacement” is perhaps a bit of a misnomer for this type of sampling scheme. Nevertheless, we use it in order to maintain consistency with [MW25]. We follow the strategy of [MW25] for sampling indices without replacement.
@@
A.5.1 Approximate Matrix-Vector Multiplication
We derive the following guarantee for the matrix-vector multiplication estimator for sampling without replacement. Since the number of samples are not fixed, we instead calculate the estimation error for a fixed expected number of samples (which is determined by the probabilities $q_{1},\dotsc,q_{r}$).
Theorem A.12 (Matrix-Vector multiplication, sampling without replacement).
@@ A.5.1 Approximate Matrix-Vector Multiplication
[…]
@@
Setting the right-hand-side of the above inequality equal to $\delta$ and solving for $\varepsilon$ gives us Eq.˜A.75. ∎
We can use the above result to derive the error bounds in terms of the function $\gamma(\bm{z})$ defined in Eq.˜2.3.
@@
where we use the fact that $q_{i}=cp_{i}$ when $q_{i}<1$ in the second line, the fact that the terms are non-negative in the third line, and the definition of $\gamma$ (see Eq.˜2.3) and $c\geq m_{0}$ in the third line. The result then follows from Theorem˜A.12. ∎
@@ A.5.2 Subspace Embedding
The result then follows from ˜A.3. ∎
@@ A.5.3 Sketched Regression
We now combine approximate matrix-vector multiplication guarantee (Corollary˜A.13) and subspace embedding guarantee (Theorem˜A.15) to obtain guarantee for the sketched regression estimator constructed by sampling without replacement.
@@
We closely follow the proof of Theorem˜A.5. Since $\bm{y}^{*}$ is the solution of $\min_{\bm{y}}\left\lVert\bm{U}\bm{y}-\bm{b}\right\rVert^{2}$, $\bm{U}\bm{y}^{*}-\bm{b}$ lies in the orthogonal complement of the range of $\bm{U}$, and therefore, $\bm{U}^{\mathsf{T}}(\bm{U}\bm{y}^{*}-\bm{b})=0$. Then, taking $\bm{z}=\bm{U}\bm{y}^{*}-\bm{b}=(\bm{I}-\bm{U}\bm{U}^{\mathsf{T}})\bm{b}$ in Corollary˜A.13, we can infer that using
@@
we have with probability exceeding $1-\delta/2$,
@@
then, with probability exceeding $1-\delta/2$,
@@
The remainder of the proof is the same as that of Theorem˜A.5. ∎
@@ Appendix B Description of past estimators
In this section we provide more details on how several existing estimators fit into the unified framework described in Section˜2.
@@ B.1 KernelSHAP
KernelSHAP makes use of a subsampled and reweighted version of the constrained regression formulation Eq.˜1.2 of the Shapley values. Specifically, denoting $\bm{Z}_{S}$ to be the $S$-th row of $\bm{Z}$, observe that
@@
The KernelSHAP estimator [LL17] is then defined as
@@
As noted by [MW25], this can be viewed as a constrained sketched regression problem
$\bm{\phi}^{\textup{KS}}=\operatornamewithlimits{argmin}_{\begin{subarray}{c}\bm{\phi}\in\mathbb{R}^{d}\\ \bm{1}^{\mathsf{T}}\bm{\phi}=v([d])-v(\emptyset)\end{subarray}}\left\lVert\bm{S}(\bm{Z}^{\prime}\bm{\phi}-\bm{b})\right\rVert^{2}$ (B.3)
Performing the same change of variables as in the proof of Theorem˜2.1 we find that
$\bm{\phi}^{\textup{KS}}$ $=\bm{Q}\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\left\lVert\bm{S}(\bm{Z}^{\prime}(\bm{Q}\bm{x}+\alpha\bm{1})-\bm{b})\right\rVert^{2}+\alpha\bm{1}$ (B.4)
$=\bm{Q}\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\left\lVert\bm{S}(\bm{Z}^{\prime}\bm{Q}\bm{x}-(\bm{b}-\alpha\bm{Z}^{\prime}\bm{1}))\right\rVert^{2}+\alpha\bm{1}$ (B.5)
$=\bm{Q}\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\left\lVert\bm{S}(\bm{U}\bm{x}-\bm{b}_{\alpha})\right\rVert^{2}+\alpha\bm{1}.$ (B.6)
@@ B.2 Unbiased KernelSHAP
In [CL20], the authors observe that the Shapley values can be expressed as
@@
Since $\bm{Q},d^{-1/2}\bm{1}]$ form an orthonormal basis for $\mathbb{R}^{d}$,
@@ B.3 LeverageSHAP
In [MW25], the authors show the typical formulation of the Shapley values Eq.˜1.2 can be rewritten as an unconstrained problem
@@ Appendix C Equivalence between Lagrangian and Change of Variable Framework
$\bm{\phi}^{\textup{R}}=\operatornamewithlimits{argmin}_{\begin{subarray}{c}\bm{\phi}\in\mathbb{R}^{d}\\ \bm{1}^{\mathsf{T}}\bm{\phi}=\alpha\end{subarray}}\left\lVert\bm{C}\bm{\phi}-\bm{y}\right\rVert^{2}.$ (C.1)
@@
$\bm{\phi}^{u}=\operatornamewithlimits{argmin}_{\begin{subarray}{c}\bm{\phi}\in\mathbb{R}^{d}\end{subarray}}\left\lVert\bm{C}\bm{\phi}-\bm{y}\right\rVert^{2}=\bm{M}^{+}\bm{g}$ (C.2)
@@
In order to solve Eq.˜C.1, the Lagrangian method writes,
@@
As discussed in Section˜A.3, an alternative method to solve the constrained least squares is using the change of variable to explicitly enforce the constraint. Specifically, we re-parameterize $\bm{\phi}$ as,
@@
$\operatornamewithlimits{argmin}_{\begin{subarray}{c}\bm{x}\end{subarray}}\left\lVert\bm{C}(\alpha\bm{1}+\bm{Q}\bm{x})-\bm{y}\right\rVert^{2}=\operatornamewithlimits{argmin}_{\begin{subarray}{c}\bm{x}\end{subarray}}\left\lVert\bm{C}\bm{Q}\bm{x}-(\bm{y}-\alpha\bm{C1})\right\rVert^{2}$ (C.6)
@@
The second term in Eq.˜C.8 can be seen as a projection of the vector $\bm{\phi}^{u}-\alpha\bm{1}$ into the span of $\bm{Q}$ (or alternatively on the null space of $\bm{1}^{\mathsf{T}}$) with the projection matrix,
@@
Thus, plugging in this Eq.˜C.11 results in
@@
thus recovering Eq.˜C.4 by noting that $\bm{1}^{\mathsf{T}}\bm{1}=d$.
Appendix D Ratio of mean squared errors
In Section˜A.2, we saw that $\gamma(\bm{z})$ and $\eta$ (see Eq.˜2.3) give upper bounds on the sample complexity of matrix-vector multiplication and regression estimators for sampling with replacement. In this section, we study the ratio of mean squared errors for different sampling strategies these estimators in the finite-sample/asymptotic regime. We find that this ratio is determined by $\gamma$ for both these estimators, as summarized below.